OpenStudy (anonymous):

in differential equations, how did the particular solution of ydx+(2x+3y)dy = 0 became xy^2+y^3=c??? The answer that I got was ln y+1/3ln((x/y)+1)=c....how did that happen?? plssss help

OpenStudy (anonymous):

will @ganeshie8 know?

OpenStudy (anonymous):

the answer I got was ln y + 1/3ln(x/y + 1) = c

ganeshie8 (ganeshie8):

did u solve it by substitution ?

OpenStudy (anonymous):

yes... I used x=uy; dx=udy+ydu

OpenStudy (anonymous):

then I arrived with this answer, lny+1/3ln((x/y) + 1) = c... but the book's answer is xy^2+y^3=c

ganeshie8 (ganeshie8):

both answers are equivalent

ganeshie8 (ganeshie8):

just simplify the log in ur answer

OpenStudy (anonymous):

but how did that happen?? can you show me how??

OpenStudy (anonymous):

when I try to simplify, i get \[\ln(y(x/y)+1)^{1/3}=c\]

OpenStudy (anonymous):

\[\ln(y(x/y+1)^{1/3})=c\]

OpenStudy (anonymous):

how did that turn into xy^2+y^3=c?? @ganeshie8

ganeshie8 (ganeshie8):

lny+1/3ln((x/y) + 1) = c multiply thru by 3 : 3lny + ln(x/y + 1) = C lny^3 + ln(x/y +1) = C ln(y^3(x/y)+1) = C ln(xy^2 + y^3) = C

OpenStudy (anonymous):

thnks alot..

ganeshie8 (ganeshie8):

yw